OpenStudy (anonymous):

anti derivative of (x^-6+1/9sqrtx)

OpenStudy (turingtest):

\[\int\limits x^{-6}+{1\over9}\sqrt xdx\]right?

OpenStudy (anonymous):

\[(x ^{-6}+\left(\begin{matrix}1 \\ \sqrt[9]{x} \end{matrix}\right)\]

OpenStudy (turingtest):

convert the radical expression to a fractional exponent and the problem is very straightforward.

OpenStudy (anonymous):

how do i do that?

OpenStudy (turingtest):

Did you forget how to write \[{1 \over \sqrt[9]x}\]as a fractional exponent? Remember \[x^{-a}={1\over x^a}\] and \[x^{a/b}=\sqrt[b]{x^a}\]That should help you out here.

OpenStudy (anonymous):

soo x^1/9?? sorry my brain is fried..

OpenStudy (turingtest):

\[x^{-1/9} \]because it is in the denominator

OpenStudy (anonymous):

to find the der tho i need to bring it up correct?

OpenStudy (turingtest):

der?

OpenStudy (anonymous):

ya i have to find the anti der of the problem

OpenStudy (turingtest):

anti derivative=integral, same thing you don't HAVE to bring it up, but it makes the answer easier to see.